If 18.1 mol of Ar gas occupies 111.5 L, how many mL would occupy 59.3 mol occupy under the same temperature and pressure? Record your answer in scientific notation using 3 significant figures.
Answer #1

Answer

365301ml

Explanation

Ideal gas equation is

PV = nRT

V = nRT/P

Temperature and pressure is constant ,so

V = n k

V1 = n1 k

V2 = n2 k

V1/V2 = n1/n2

V1 = 111.5L

V2 = ?

n1 = 18.1mol

n2 = 59.3mol

V2 = V1×n2/n1

= 111.5L × 59.3mol /18.1mol

= 365.301L

= 365301ml